Matthew owns a specialty coffee shop.He wants to create a special mix using two coffees,one priced at $6.40 per pound and the other priced at $7.28 per pound. How many pounds of the $7.28 coffee should he mix with 9 pounds of the $6.40 coffee to sell the mixture for $6.95 per pound?
Found 2 solutions by Paul, wuwei96815:
Answer by Paul(988)

728x+640(9-x)=695(9)
728x+5760-640x=6255
88x=495
x=5.625
9-5.625=3.375
Hence, for $7.28 there should be 5.625 of solution added and for $6.40 there should be 3.375 of solution added.

Answer by wuwei96815(245)

The hardest part of word problems is translating. What are they really asking for? Once you can figure that out, the math is usally relatively easy.

Let x=the unknown quantity of coffee at $7.28/pound.

7.28(x) + 6.40(9) = (x+9)6.95
7.28x + 57.60 = 6.95x + 62.55
7.28x-6.95x = 62.55-57.60
0.33x = 4.95
x = 15 pounds of coffee at $7.28/pound

I hope that is right?